In a titration of 42.91 mL of 0.3566 M ammonia with 0.3566 M aqueous nitric acid, what is the pH of the solution when 42.91 mL of the acid have been added?
Answer #1

Given:-

molarity of ammonia (NH3) = 0.3566 M = 0.3566 mole / L

volume of ammonia (NH3) used = 42.91 mL  = 42.91/ 1000 = 0.04291 L

molarity of nitric acid (HNO3) = 0.3566 M   = 0.3566 mole / L

volume of nitric acid (HNO3) used = 42.91 mL = 42.91/ 1000 = 0.04291 L

pH of resultant solution = ?

As we know that

molarity of compound (M) = no. of moles of compound (n) / volume of solution (L)

no. of moles of compound (n) = molarity of compound (M) \times volume of solution (L)

Since we know that nitric acid (HNO3) is a strong acid therefore

no. of moles of nitric acid (HNO3) (nHNO3) = molarity of nitric acid (HNO3) (M) \times volume of solution (L)

no. of moles of nitric acid (HNO3) (nHNO3) = 0.3566 mole / L  \times 0.04291 L

no. of moles of nitric acid (HNO3) (nHNO3) = 0.015302 moles

similarly

no. of moles of ammonia (NH3) (nNH3) = molarity of ammonia (NH3) (M) \times volume of solution (L)

no. of moles of ammonia (NH3) (nNH3) =  0.3566 mole / L  \times 0.04291 L

no. of moles of ammonia (NH3) (nNH3) = 0.015302 moles

Since we know that  when weak base such as ammonia (NH3) being titrated with a strong acid such as nitric acid (HNO3), the neutralization reaction occur and that will produce the conjugate acid (NH4+) of the weak base (NH3), which will then react with water to produce hydronium ions, H3O+ and this hydronium ion determine the pH of the solution.

NH3(aq) + HNO3(aq) \rightarrow   NH4NO3(aq)

above reaction can be written as follows:-

NH3(aq) + HNO3(aq) \rightarrow NH4+(aq) + NO3-(aq)

0.015302 moles   0.015302 moles   0.015302 moles

So

No. of moles of NH4+ (nNH4+) = 0.015302 moles

Also know that the total volume of solution

Total volume of solution (Vtotal) = volume of ammonia (NH3) used + volume of nitric acid (HNO3) used

Total volume of solution (Vtotal) = 0.04291 L + 0.04291 L

Total volume of solution (Vtotal) = 0.08582 L

therefore

molar concentration of NH4+ i.e [NH4+] = No. of moles of NH4+ (nNH4+) /  Total volume of solution (Vtotal)

molar concentration of NH4+ i.e [NH4+] = 0.015302 moles /  0.08582 L

molar concentration of NH4+ i.e [NH4+] = 0.1783 mole /L = 0.1783 M

So we know that molar concentration of NH4+ would determine the molar concentration of hydronium ions, H3O+

NH4+(aq) + H2O(l) \rightleftharpoons NH3(aq)    + H3O+(aq)

Initial 0.1783 0 0

Change -x +x + x

Equiulibrium  0.1783 - x     +x + x

Now

acid dissociation constant of NH4+ (Ka) = [NH3][H3O+] / [NH4+]

acid dissociation constant of NH4+ (Ka) = x \times x  / (0.1783 - x )

acid dissociation constant of NH4+ (Ka) = x2 / (0.1783 - x )

since

0.1783 - x  \simeq  0.1783

so above equation becomes

acid dissociation constant of NH4+ (Ka) = x2 / 0.1783 --------------------------------(1)

Also we know that

NH3(aq) + H2O(l) \rightleftharpoons   NH4+(aq) + OH-(aq)

base dissociation constant of NH3 (Kb) = [NH4+][OH-] / [NH3 ]

1.8 \times 10-5 = [NH4+][OH-] / [NH3 ]

i.e

base dissociation constant of NH3 (Kb) = 1.8 \times 10-5

since we know that

KaKb = Kw

Ka = Kw / Kb

therefore

acid dissociation constant of NH4+ (Ka) = Kw / base dissociation constant of NH3 (Kb)

acid dissociation constant of NH4+ (Ka) = 10-14 / 1.8 \times 10-5

acid dissociation constant of NH4+ (Ka) = 5.6 \times 10-10 ----------------------------------(2)

form equation no. 1 and 2 we get

5.6 \times 10-10 = x2 / 0.1783

x2 = 0.1783 \times 5.6 \times 10-10

x2 = 0.99848 \times 10-10

x = \sqrt{} 0.99848 \times 10-10

x =  0.999 \times 10-5

therefore

[H3O+] = 0.999 \times 10-5

As we know that

pH = - log[H3O+]

pH = - log[0.999 \times 10-5]

pH = - [ log(0.999) + ( - 5log(10) ]

pH = 5log(10) - log(0.999)

pH = 5  - (- 4.36)

pH = 5  + 4.36

pH = 9.36 (i.e answer)